find the composition
f (x) = -2x-2 and g (x)= x^2+x
find F of g (5)

Answer by nyc_function(2741)

We need to find g(5) first.
g(x) = x^2 + x
g(5) = (5)^2 + 5
g(5) = 25 + 5
g(5) = 30
We now find f(g(5)). What is g(5)?
We just learned that it is 30.
In other words, g(5) = 30 and vice-versa.
f(g(5)) = -2(30) - 2
f(g(5)) = -60 - 2
f(g(5)) = -62
Did you follow?